‘X’ is molecular substance containing 21 atoms of carbon per molecule. The weight % of C in ‘X’ is 50.4%.What is molecular weight of X in amu.?

Grade:11

1 Answers

Sheetal kumari
6 years ago
consider the moecular substance be XC21 containing 21 atoms of carbon.As we know that Mass % of an element = mass of the element in the compound /molar mass of the compound.mass of carbon in this compound = 12(atomic mass )×21(atoms of carbon ).=252according to question we have 50.4=252×100 /molar mass Molar mass =500gPlease kindly approved my answer